OpenStudy (i_love_my_nieces):

The ordered pairs (2, -11) and (5, -20) are solutions to which of the following equations? y = 3x - 5 y = -3x + 5 y = 3x + 5 y = -3x - 5

OpenStudy (mrnood):

This is exactly the same as you had help on before. Just take the (x,y) values from the brackets and put tem into the equation. If the result is 'correct' or true then try the other brackets in the same equation. Only ONE of the above equations will give oyu a true answer for both pairs: For example (2,-11) Equation is y= 3x-5 put y= -11 put x = 2 equation becomes -11=3(2)-5 But 3(2)-5 = 1 so this is NOT true so the first one I have done for you TRY the other 3 I'm sure you can do it.....

OpenStudy (i_love_my_nieces):

@ganeshie8

OpenStudy (mrnood):

So - you are not even going to TRY?